# Circular geodesic orbits in the equatorial plane of a charged rotating disc of dust

### v1

@@ Abstract
Equatorial circular geodesic orbits of neutral test particles in the exterior spacetime of the charged rotating disc of dust are analyzed in dependence of its specific charge and a relativity parameter. The charged rotating disc of dust is an axisymmetric, stationary solution of the Einstein-Maxwell equations in terms of a post-Newtonian expansion. In particular, photon, marginally bound and marginally stable orbits are discussed. It turns out that general formulae in closed form for angular velocity, specific energy and specific angular momentum of the test particles can be derived, which hold for any (exterior) asymptotically flat, axisymmetric, stationary and reflection symmetric (electro-)vacuum spacetime. Furthermore, circular motion in the exterior spacetime of the charged rotating disc of dust is qualitatively very similar to that around a Kerr-Newman black hole for sufficiently large radii, but differs strongly in the respective closer vicinity.
@@ I Introduction
Meinel and Kleinwächter [5] (see also [6]) and Ansorg [7] analyzed (timelike) geodesic motion in the gravitational field of the uncharged rotating disc of dust [8, 9, 6].
In the present paper circular orbits of neutral test particles in the equatorial plane of an isolated, axisymmetric, stationary, reflection symmetric, charged, rotating body are investigated first. Remarkably, even without knowing the explicit solution of the rotating body, energy, angular momentum and angular velocity of the test particles can be specified in closed form solely as functions of the metric and their radial derivatives.
The general examination of equatorial circular orbits of neutral test particles is followed by a specialization on the spacetime of the charged rotating disc of dust. This disc is rigidly rotating, has constant specific charge and is an axisymmetric, stationary solution of the Einstein-Maxwell equations expressed in terms of a post-Newtonian expansion up to tenth order [10, 11]. A detailed discussion of equatorial circular orbits in the spacetime of the charged rotating disc of dust is provided.
@@ II Charged rotating disc of dust
Note that the relativity parameter can equivalently be expressed in in terms of the redshift $z_{c}$ of a photon traveling from the center of the disc to spatial infinity:
@@
The specific charge $\epsilon$, furthermore, directly influences the disc’s rotation speed. As is evident from the Newtonian limit (where each dust particle within the disc is in an equilibrium of gravitational, electric and centrifugal force), disc configurations with $\epsilon=0$ have a maximum angular velocity and those with $\epsilon=1$ are static.
@@ III Circular geodesic orbits in a general spacetime
We first consider the exterior spacetime of an isolated, axisymmetric, stationary, reflection symmetric, charged, rotating body. These qualifications, besides a non-zero charge, are expected to be fulfilled for most astrophysical bodies. (Note that the special cases of zero charge and vanishing rotation are included.)
@@
where $\dot{x}^{a}=\frac{\mathrm{d}x^{a}}{\mathrm{d}\tau}$. The resulting Euler-Lagrange equation is equivalent to the geodesic equation, i.e. the extremal path between two points is also straightest possible path.
@@
Differentiation of eq. 23 with respect to the particles proper time $\tau$ leads to the equation of motion:
@@
Inserting the metric functions of the uncharged rotating disc of dust at the rim into the above formulae (of the first or second approach) reveals the same expressions for $\Omega_{1/2}$, $L_{1/2}$ and $E_{1/2}$ (with 1/2=+/1/2=+/-1 / 2 = + / -) as in [5]. Also the corresponding results for equatorial circular geodesic orbits in the (exterior) Kerr [1] and Kerr-Newman spacetime [2] can be obtained from the above formulae (using either approach). [^8] Note that $\rho\geq 0$.
@@
IV Circular geodesic orbits in the spacetime of the charged rotating disc of dust
Now circular geodesic orbits in the equatorial plane ($\eta=0$) of the charged rotating disc of dust are to be discussed. By evaluating eqs. 15, 16 and 17 or eqs. 32, 29, 30 and 31 with the metric of the disc, using eqs. 4 and 5, the angular velocity $\Omega_{1/2}=\Omega_{+/-}$, the specific angular momentum $L_{1/2}=L_{+/-}$ and the specific energy $E_{1/2}=E_{+/-}$ can be expressed in terms of a post-Newtonian expansion, see appendix A. Here “$+$” denotes direct and “--” retrograde orbits. The angular velocity of the disc, $\Omega$, is non-negative.
@@ IV Circular geodesic orbits in the spacetime of the charged rotating disc of dust
It was further confirmed that the simultaneous limit $\epsilon\to 0$, $\rho\to\rho_{0}$ leads to the same results for $\Omega_{\pm}$ and $L_{\pm}$, $E_{\pm}$, in terms of a series expansion in $g$, as in [5]. Especially noteworthy is the fact that $\Omega_{+}=\Omega$ in this limiting case. The reason is that the $\rho$-component of the geodesic equation of a test particle, eq. 12, also represents a boundary condition of the uncharged rotating disc of dust (see [6]).
@@
The convergence behavior of $R_{0}\Omega_{\pm}$ and $\frac{L_{\pm}}{R_{0}}$ is computed analogously. Generally, the quantities corresponding to prograde circular motion exhibit a significantly better convergence behavior than those of retrograde circular motion. It should be noted that $\Omega_{\pm}$ and $L_{\pm}$ are only exact up to ninth order.

### v2

@@ Abstract
Equatorial circular geodesic orbits of neutral test particles in the exterior spacetime of a charged rotating disc of dust are analyzed in dependence of its specific charge and a relativity parameter. The charged rotating disc of dust is an axisymmetric, stationary solution of the Einstein-Maxwell equations in terms of a post-Newtonian expansion. In particular, photon, marginally bound and marginally stable orbits are discussed. It turns out that general formulae in closed form for angular velocity, specific energy and specific angular momentum of the test particles can be derived, which hold for any (exterior) asymptotically flat, axisymmetric, stationary and reflection symmetric (electro-)vacuum spacetime. Furthermore, circular motion in the exterior spacetime of the charged rotating disc of dust is qualitatively very similar to that around a Kerr-Newman black hole for sufficiently large radii, but differs strongly in the respective closer vicinity.
@@ I Introduction
Meinel and Kleinwächter [5] (see also [6]) and Ansorg [7] analyzed (timelike) geodesic motion in the gravitational field of an uncharged rotating disc of dust [8, 9, 6] (which may serve as a basic model for disc galaxies).
In the present paper circular orbits of neutral test particles in the equatorial plane of an isolated, axisymmetric, stationary, reflection symmetric, charged, rotating body are investigated first. This class includes black holes (in particular the Kerr-Newman solution), rotating disc configurations (such as rigidly rotating discs of dust with or without charge) and generally (idealized, uncharged) astrophysical objects. Remarkably, even without knowing the explicit solution of the rotating body, energy, angular momentum and angular velocity of the test particles can be specified in closed form solely as functions of the metric and their radial derivatives.
The general examination of equatorial circular orbits of neutral test particles is followed by a specialization on the spacetime of a charged rotating disc of dust. This disc is rigidly rotating, has constant specific charge and is an axisymmetric, stationary solution of the Einstein-Maxwell equations expressed in terms of a post-Newtonian expansion up to tenth order [10, 11]. A detailed discussion of equatorial circular orbits in the spacetime of the charged rotating disc of dust is provided.
@@ II Charged rotating disc of dust
Note that the relativity parameter can equivalently be expressed in terms of the redshift $z_{c}$ of a photon traveling from the center of the disc to spatial infinity:
@@
The specific charge $\epsilon$, furthermore, directly influences the disc’s rotation speed. As is evident from the Newtonian limit (where each dust particle within the disc is in an equilibrium of gravitational, electric and centrifugal force), disc configurations with $\epsilon=0$ have a maximum angular velocity and those with $\epsilon=1$ are static. [^3] A detailed investigation of the dependence of the gravitational mass $M$, the angular momentum $J$, the electric charge $Q$ and the magnetic dipole moment $D$ (each with an appropriate normalization), as well as higher (normalized) multipole moments, on the parameters $g$ and $\epsilon$ can be found in [16].
@@ III Circular geodesic orbits in a general spacetime
We first consider the exterior spacetime of an isolated, axisymmetric, stationary, reflection symmetric, charged, rotating body. These qualifications, besides a non-zero charge, are expected to be fulfilled for most astrophysical bodies. [^4] Axisymmetry, stationarity and reflection symmetry should be fulfilled to a good approximation for most astrophysical objects. (Note that the special cases of zero charge and vanishing rotation are included.)
@@
where $\dot{x}^{a}=\frac{\mathrm{d}x^{a}}{\mathrm{d}\tau}$. The resulting Euler-Lagrange equation is equivalent to the geodesic equation, i.e. the extremal path between two points is also the straightest possible path.
@@
Differentiation of eq. 23 with respect to the particle proper time $\tau$ leads to the equation of motion:
@@
Inserting the metric functions of the uncharged rotating disc of dust [8, 9, 6] at the rim into the above formulae (of the first or second approach) reveals the same expressions for $\Omega_{1/2}$, $L_{1/2}$ and $E_{1/2}$ (with 1/2=+/1/2=+/-1 / 2 = + / -) as in [5]. Also the corresponding results for equatorial circular geodesic orbits in the (exterior) Kerr [1] and Kerr-Newman spacetime [2] can be obtained from the above formulae (using either approach). [^10] Note that $\rho\geq 0$.
IV Circular geodesic orbits in the spacetime of a charged rotating disc of dust
@@
Now circular geodesic orbits in the equatorial plane ($\eta=0$) of the charged rotating disc of dust, described in section II, are to be discussed. By evaluating eqs. 15, 16 and 17 or eqs. 32, 29, 30 and 31 with the metric of the disc, using eqs. 4 and 5, the angular velocity $\Omega_{1/2}=\Omega_{+/-}$, the specific angular momentum $L_{1/2}=L_{+/-}$ and the specific energy $E_{1/2}=E_{+/-}$ can be expressed in terms of a post-Newtonian expansion, see appendix A. Here “$+$” denotes direct and “--” retrograde orbits. The angular velocity of the disc, $\Omega$, is non-negative.
@@
It was further confirmed that the simultaneous limit $\epsilon\to 0$, $\rho\to\rho_{0}$ leads to the same results for $\Omega_{\pm}$, $L_{\pm}$ and $E_{\pm}$, in terms of a series expansion in $g$, as in [5]. Especially noteworthy is the fact that $\Omega_{+}=\Omega$ in this limiting case. The reason is that the $\rho$-component of the geodesic equation of a test particle, eq. 12, also represents a boundary condition of the uncharged rotating disc of dust (see [6]).
@@
The convergence behavior of $R_{0}\Omega_{\pm}$ and $\frac{L_{\pm}}{R_{0}}$ is computed analogously. Generally, the quantities corresponding to prograde circular motion exhibit a significantly better convergence behavior than those of retrograde circular motion. It should be noted that $\Omega_{\pm}$ and $L_{\pm}$ are only exact up to ninth order. [^16] Expanding a quantity that itself involves series expansions up to a certain order does not necessarily lead to a series expansion that is exact up to this order.